OpenStudy (anonymous):

Statistics question: find the probability that there are some left handed people in a group of 5. Assuming that the probability of being left handed is .1

OpenStudy (shane_b):

I assume this is as simple as:\[\frac{.1}{1}=\frac{10}{100}=\frac{0.5}{5}\]So about half a person would be left handed out of 5 people :P

OpenStudy (anonymous):

p=0.1 , n=5 So, P(x>=1)=1-P(x=0)=1-5C0 * p^0 * (1-p)^5=0.41

OpenStudy (anonymous):

Thus, the probability that there are some left handed people in a group of 5 is 0.41

OpenStudy (anonymous):

half a person walks around left-handed... but his other half is right-handed. oh wait there's a word for that - ambidextrous :)